# Text-to-Vision Generation with Scene Graph Programming

### v2

@@ 9 Conclusion
We present Generate Any Scene, a system leveraging scene graph programming to generate diverse and compositional synthetic captions for Text-to-Vision generation tasks. It extends beyond existing real-world caption datasets to include imaginary scenes and even implausible scenarios. To demonstrate the effectiveness of Generate Any Scene, we explore three applications: (1) self-improvement by iteratively optimizing models, (2) distillation of proprietary model strengths into open-source models, and (3) robust content moderation with diverse synthetic data. Generate Any Scene highlights the importance of synthetic data in evaluating and improving Text-to-Vision generation, and addresses the need to systematically define and scalably produce the space of visual scenes.
Acknowledgement
This project was partially supported by an OpenAI “Superalignment Fast” grant.
Supplementary Material
Appendix A More Analysis with Generate Any Scene
With Generate Any Scene, we can generate infinitely diverse and highly controllable prompts. Using Generate Any Scene, we conduct several analyses to provide insights into the performance of today’s Text-to-Vision generation models.
A.1 Performance analysis across caption properties
Perplexity. (Figure 8)
Commonsense. (Figure 9)
Commonsense is an inherent property of text. We utilize the Vera Score [40], a metric generated by a fine-tuned LLM to evaluate the text’s commonsense level.
Element Numbers (Complexity of Scene Graph). (Figure 10)
A.2 Analysis on different metrics
Clip Score isn’t a universal metric.
Clip Score is one of the most widely used metrics in Text-to-Vision generation for evaluating the alignment between visual content and text. However, our analysis reveals that Clip Score is not a perfect metric and displays some unusual trends. For instance, as shown in Figures 8, 9, and 10, we compute the perplexity across 10k prompts used in our study, where higher perplexity indicates more unpredictable or disorganized text. Interestingly, unlike other metrics, Clip Score decreases as perplexity lowers, suggesting that Clip Score tends to favor more disorganized text. This behavior is counterintuitive and highlights the potential limitations of using Clip Score as a robust alignment metric.
Limitations of human preference-based metrics.
We use two metrics fine-tuned using human preference data: Pick Score and ImageReward Score. However, we found that these metrics exhibit a strong bias toward the data on which they were fine-tuned. For instance, as shown in Table 5, Pick Score assigns similar scores across all models, failing to provide significant differentiation or meaningful insights into model performance. In contrast, ImageReward Score demonstrates clearer preferences, favoring models such as DaLL-E 3 and Playground v2.5, which incorporated human-alignment techniques during their training. However, this metric shows a significant drawback: it assigns disproportionately large negative scores to models like Stable Diffusion v2-1, indicating a potential over-sensitivity to alignment mismatches. Such behavior highlights the limitations of these metrics in providing fair and unbiased evaluations across diverse model architectures.
VQA Score and TIFA Score are relative reliable metrics.
A.3 Fairness analysis
Gender.
Figure 11: Average performance scores of all models across different genders were evaluated using various metrics.
Race.
There are also performance gaps in different races. From Figure 12, we found that ”white (person)” and ”black (person)” perform better than ”asian (person)”, ”Indian (amerindian)”, and ”Latin American”.
@@ A.3 Fairness analysis
Appendix B Correlation of Generate Any Scene with other Text-to-Vision generation benchmarks
The Generate Any Scene benchmark uniquely relies entirely on synthetic captions to evaluate models. To assess the transferability of these synthetic captions, we analyzed the consistency in model rankings across different benchmarks [76, 34, 75]. Specifically, we identified the overlap of models evaluated by two benchmarks and computed the Spearman correlation coefficient between their rankings.
As shown in the figure 13, Generate Any Scene demonstrates a strong correlation with other benchmarks, such as Conceptmix [76] and GenAI Bench [34], indicating the robustness and reliability of Generate Any Scene’s synthetic caption-based evaluations. This suggests that the synthetic captions generated by Generate Any Scene can effectively reflect model performance trends, aligning closely with those observed in benchmarks using real-world captions or alternative evaluation methods.
Appendix C Details of Taxonomy of Visual Concepts
Our metadata is further organized using a well-defined taxonomy, enhancing the ability to generate controllable prompts. This hierarchical taxonomy not only facilitates the creation of diverse scene graphs, but also enables fine-grained and systematic model evaluation.
Objects.
To enhance the comprehensiveness and taxonomy of object data, we leverage noun synsets and the structure of WordNet [45]. In WordNet, a physical object is defined as ”a tangible and visible entity; an entity that can cast a shadow.” Following this definition, we designate the physical object as the root node, constructing a hierarchical tree with all 28,787 hyponyms under this category as the set of objects in our model.
Attributes.
The attributes of a scene graph represent properties or characteristics associated with each object. We classify these attributes into nine primary categories. For color, we aggregate 677 unique entries sourced from Wikipedia [72]. The material category comprises 76 types, referenced from several public datasets [64, 4, 79]. The texture category includes 42 kinds from the Describable Textures Dataset [14], while the architectural style encompasses 25 distinct styles [78]. Additionally, we collect 85 states, 41 shapes, and 24 sizes. For human descriptors, we compile 59 terms across subcategories, including body type and height. Finally, we collect 465 common adjectives covering general characteristics of objects to enhance the descriptive richness of our scene graphs.
Relationships.
We leverage the Robin dataset [49] as the foundation for relationship metadata, encompassing six key categories: spatial, functional, interactional, social, emotional, and symbolic. With 10,492 relationships, the dataset provides a comprehensive and systematic repository that supports modeling diverse and complex object interactions. Its extensive coverage captures both tangible and abstract connections, forming a robust framework for accurate scene graph representation.
Scene Attributes.
In Text-to-Vision generation tasks, people mainly focus on creating realistic images and art from a text description [58, 55, 5]. For artistic styles, we define scene attributes using 76 renowned artists, 41 genres, and 126 painting styles from WikiArt [59], along with 29 common painting techniques. For realistic imagery, we construct camera settings attributes across 6 categories: camera models, focal lengths, perspectives, apertures, depths of field, and shot scales. The camera models are sourced from the 1000 Cameras Dataset [15], while the remaining categories are constructed based on photography knowledge and common prompts in Text-to-Vision generation tasks [7, 70]. To control scene settings, we categorize location, weather and lighting attributes, using 430 diverse locations from Places365 [41], alongside 76 weathers and 57 lighting conditions. For video generation, we introduce attributes that describe dynamic elements. These include 12 types of camera rig, 30 distinct camera movements, 15 video editing styles, and 27 temporal spans. The comprehensive scene attributes that we construct allow for the detailed and programmatic Text-to-Vision generation generation.
Appendix D Details of Overall Performance (Section 3)
D.1 Detailed experiment settings
@@
    For Text-to-Image generation, we select a range of open-source models, including those utilizing UNet backbones, such as DeepFloyd IF [1], Stable Diffusion v2-1 [56], SDXL [52], Playground v2.5 [35], and Wuerstchen v2 [51], as well as models with DiT backbones, including Stable Diffusion 3 Medium [16], PixArt-$\alpha$ [11], PixArt-$\Sigma$ [12], FLUX.1-schnell [33], FLUX.1-dev [33], and FLUX 1. Closed-source models, such as DaLL-E 3 [5] and FLUX1.1 PRO [33], are also assessed to ensure a comprehensive comparison. All models are evaluated at a resolution of 1024 × 1024 pixels.
@@
    For Text-to-Video generation, we select nine open-source models: ModelScope [67], ZeroScope [61], Text2Video-Zero [30], CogVideoX-2B [80], VideoCrafter2 [10], AnimateLCM [65], AnimateDiff [19], FreeInit [74], and Open-Sora 1.2 [84]. We standardize the frame length to 16 across all video models for fair comparisons.
@@
    For Text-to-3D generation, we evaluate five recently proposed models: SJC [66], DreamFusion [53], Magic3D [38], Latent-NeRF [44], and ProlificDreamer [69]. We employ the implementation and configurations provided by ThreeStudio [18] and generate videos by rendering from 120 viewpoints. To accelerate inference, we omit the refinement stage. For Magic3D and DreamFusion, we respectively use DeepFloyd IF and Stable Diffusion v2-1 as their 2D backbones.
Metrics.
Across all Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation, we employ five widely used Text-to-Vision generation metrics to comprehensively assess model performance:
@@
We also use metrics in VBench [24] to evaluate Text-to-Video generation models on fine-grained dimensions, such as consistency and dynamics, providing detailed insights into video performance.
@@
This sampling approach optimizes inference speed without affecting score accuracy [39].
D.2 Detailed overall results
We evaluate Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation models on Generate Any Scene. The detailed results of each model and each metric are shown in  Tabs. 5, 6, 7 and 8
@@ D.2 Detailed overall results
Playground v2.5 [35] 0.2581 0.2132 0.5734 0.2569 0.2919
Stable Diffusion v2-1 [56] 0.2453 0.1988 0.5282 0.2310 -0.9760
SDXL [52] 0.2614 0.2046 0.5328 0.2361 -0.3463
Wuerstchen v2 [51] 0.2448 0.2022 0.5352 0.2239 -0.3339
DeepFloyd IF XL [1] 0.2396 0.1935 0.5397 0.2171 -0.8687
Stable Diffusion 3 Medium [16] 0.2527 0.2027 0.5579 0.2693 -0.0557
PixArt-$\alpha$ [11] 0.2363 0.2050 0.6049 0.2593 0.1149
PixArt-$\Sigma$ [12] 0.2390 0.2068 0.6109 0.2683 0.0425
FLUX.1-dev [33] 0.2341 0.2060 0.5561 0.2295 0.1588
FLUX.1-schnell [33] 0.2542 0.2047 0.6132 0.2833 0.1251
FLUX1.1 PRO [33] 0.2315 0.2065 0.5744 0.2454 -0.0361
Dalle-3 [5] 0.2518 0.2006 0.6871 0.4249 0.3464
Table 5: Overall performance of Text-to-Image generation models over 10K Generate Any Scene prompts.
Model clip score pick score image reward score VQA score TiFA score
VideoCraft2 [10] 0.2398 0.1976 -0.4202 0.5018 0.2466
AnimateDiff [19] 0.2610 0.1959 -0.7301 0.5255 0.2208
Open-Sora 1.2 [84] 0.2259 0.1928 -0.6277 0.5519 0.2414
FreeInit [74] 0.2579 0.1950 -0.9335 0.5123 0.2047
ModelScope [67] 0.2041 0.1886 -1.9172 0.3840 0.1219
Text2Video-Zero [30] 0.2539 0.1933 -1.2050 0.4753 0.1952
AnimateLCM [65] 0.2450 0.1987 -0.5754 0.4816 0.2176
CogVideoX-2B [80] 0.2038 0.1901 -1.2301 0.4585 0.1997
ZeroScope [61] 0.2289 0.1933 -1.1599 0.4892 0.2388
Table 6: Overall performance of Text-to-Video generation models over 10k Generate Any Scene prompts.
Model subject consistency background consistency motion smoothness dynamic degree aesthetic quality imaging quality
@@
Open-Sora 1.2 0.9964 0.9907 0.9973 0.0044 0.5235 0.6648
FreeInit 0.9581 0.9571 0.9752 0.4440 0.5200 0.5456
ModelScope 0.9795 0.9831 0.9803 0.1281 0.3993 0.6494
Text2Video-Zero 0.8471 0.9030 0.8301 0.9999 0.4889 0.7018
AnimateLCM 0.9883 0.9802 0.9887 0.0612 0.6323 0.6977
@@
Table 7: Overall performance of Text-to-Image generation models over 10k Generate Any Scene prompts with VBench metrics.
Model clip score pick score vqa score tifa score image reward score
ProlificDreamer [69] 0.2125 0.1940 0.5411 0.2704 -1.2774
Latent-NeRF [44] 0.2115 0.1910 0.4767 0.2216 -1.5311
DreamFusion-sd [53] 0.1961 0.1906 0.4421 0.1657 -1.5582
Magic3D-sd [38] 0.1947 0.1903 0.4193 0.1537 -1.6327
SJC [66] 0.2191 0.1915 0.5015 0.2563 -1.4370
DreamFusion-IF [53] 0.1828 0.1857 0.3872 0.1416 -1.9353
Magic3D-IF [38] 0.1919 0.1866 0.4039 0.1537 -1.8465
@@
Table 8: Overall performance of Text-to-3D generation models over 10k Generate Any Scene prompts.
D.3 Case study: Pairwise fine-grained model comparison

### v3

@@ 9 Conclusion
We present Generate Any Scene, a system leveraging scene graph programming to generate diverse and compositional synthetic captions for Text-to-Vision generation tasks. It extends beyond existing real-world caption datasets to include comprehensive scenes and even implausible scenarios. To demonstrate the effectiveness of Generate Any Scene, we explore four applications: (1) self-improvement by iteratively optimizing models, (2) distillation of proprietary model strengths into open-source models, (3) a scene-graph-based efficient reward model within the GRPO, and (4) robust content moderation with diverse synthetic data. Generate Any Scene highlights the importance of synthetic data in improving Text-to-Vision generation, and addresses the need to systematically define and scalably produce the space of visual scenes.
@@
Appendix A Evaluating Text-to-Vision generation models with Generate Any Scene
A.1 Experiment Settings
Models. We conduct experiments on 12 Text-to-image models (at StabilityAI, 2023; Podell et al., 2023; Rombach et al., 2022b; Li et al., 2024b; Pernias et al., 2023; Esser et al., 2024; Chen et al., 2023a; 2024b; Labs, 2024; Betker et al., 2023), 9 Text-to-Video models (Wang et al., 2023b; Sterling, 2023; Khachatryan et al., 2023; Guo et al., 2023b; Wang et al., 2024a; Wu et al., 2025; Zheng et al., 2024; Yang et al., 2024; Chen et al., 2024a), and 5 Text-to-3D models (Poole et al., 2022; Lin et al., 2023; Wang et al., 2023a; 2024c; Metzer et al., 2023).
@@ A.1 Experiment Settings
    For Text-to-Image generation, we select a range of open-source models, including those utilizing UNet backbones, such as DeepFloyd IF (at StabilityAI, 2023), SDv2.1 (Rombach et al., 2022b), SDXL (Podell et al., 2023), Playground v2.5 (Li et al., 2024b), and Wuerstchen v2 (Pernias et al., 2023), as well as models with DiT backbones, including SD3 Medium (Esser et al., 2024), PixArt-$\alpha$ (Chen et al., 2023a), PixArt-$\Sigma$ (Chen et al., 2024b), FLUX.1-schnell (Labs, 2024), FLUX.1-dev (Labs, 2024), and FLUX 1. Closed-source models, such as DaLL-E 3 (Betker et al., 2023) and FLUX1.1 PRO (Labs, 2024), are also assessed to ensure a comprehensive comparison. All models are evaluated at a resolution of 1024 × 1024 pixels.
@@
    For Text-to-Video generation, we select nine open-source models: ModelScope (Wang et al., 2023b), ZeroScope (Sterling, 2023), Text2Video-Zero (Khachatryan et al., 2023), CogVideoX-2B (Yang et al., 2024), VideoCrafter2 (Chen et al., 2024a), AnimateLCM (Wang et al., 2024a), AnimateDiff (Guo et al., 2023b), FreeInit (Wu et al., 2025), and Open-Sora 1.2 (Zheng et al., 2024). We standardize the frame length to 16 across all video models for fair comparisons.
@@
    For Text-to-3D generation, we evaluate five recently proposed models: SJC (Wang et al., 2023a), DreamFusion (Poole et al., 2022), Magic3D (Lin et al., 2023), Latent-NeRF (Metzer et al., 2023), and ProlificDreamer (Wang et al., 2024c). We employ the implementation and configurations provided by ThreeStudio (Guo et al., 2023a) and generate videos by rendering from 120 viewpoints. To accelerate inference, we omit the refinement stage. For Magic3D and DreamFusion, we respectively use DeepFloyd IF and SDv2.1 as their 2D backbones.
Metrics. Across all Text-to-Vision generation tasks, we use Clip Score (Chakrabarty et al., 2023) (semantic similarity), VQA Score (Lin et al., 2024) (faithfulness), TIFA Score (Hu et al., 2023; Cho et al., 2023) (faithfulness), Pick Score (Kirstain et al., 2023) (human preference), and ImageReward Score (Xu et al., 2023) (human preference) as general metrics:
@@
We also use metrics in VBench (Huang et al., 2024) to evaluate Text-to-Video generation models on fine-grained dimensions, such as consistency and dynamics, providing detailed insights into video performance.
@@
This sampling approach optimizes inference speed without affecting score accuracy (Lin et al., 2024).
Synthetic captions. We evaluate our Text-to-Image generation and Text-to-Video generation models on 10K randomly generated captions, with scene graph complexity ranging from 3 to 12 and scene attributes from 0 to 5, using unrestricted metadata. The captions exhibit an average graph degree of 1.15, with values spanning from 0.0 to 0.8. The mean number of connected components per scene graph is 3.51, ranging from 1 to 11. For Text-to-3D generation models, due to their limitations in handling complex captions and time-intensive generation, we restrict scene graph complexity to 1-3, scene attributes to 0-2, and evaluate on 1K captions.
A.2 Overall results
We evaluate Text-to-Image generation, Text-to-Video generation, and Text-to-3D generation models on Generate Any Scene.
Table 3: Overall performance of Text-to-Image generation models over 10K Generate Any Scene captions. Evaluated on a 1K caption subset due to inference cost constraints.
@@ A.2 Overall results
Playground v2.5 (Li et al., 2024b) 0.2581 0.2132 0.5734 0.2569 0.2919
Stable Diffusion v2-1 (Rombach et al., 2022b) 0.2453 0.1988 0.5282 0.2310 -0.9760
SDXL (Podell et al., 2023) 0.2614 0.2046 0.5328 0.2361 -0.3463
Wuerstchen v2 (Pernias et al., 2023) 0.2448 0.2022 0.5352 0.2239 -0.3339
DeepFloyd IF XL (at StabilityAI, 2023) 0.2396 0.1935 0.5397 0.2171 -0.8687
Stable Diffusion 3 Medium (Esser et al., 2024) 0.2527 0.2027 0.5579 0.2693 -0.0557
PixArt-$\alpha$ (Chen et al., 2023a) 0.2363 0.2050 0.6049 0.2593 0.1149
PixArt-$\Sigma$ (Chen et al., 2024b) 0.2390 0.2068 0.6109 0.2683 0.0425
FLUX.1-dev (Labs, 2024) 0.2341 0.2060 0.5561 0.2295 0.1588
FLUX.1-schnell (Labs, 2024) 0.2542 0.2047 0.6132 0.2833 0.1251
FLUX1.1 PRO (Labs, 2024) 0.2315 0.2065 0.5744 0.2454 -0.0361
Dalle-3 (Betker et al., 2023) 0.2518 0.2006 0.6871 0.4249 0.3464
Text-to-Image generation results. (Figure 7, Table 3)
  1. 1.
  2. 2.
  3. 3.
    The closed-source model DaLL-E 3 maintains a significant lead in VQA Score, TIFA Score, and ImageReward Score, demonstrating strong faithfulness and alignment with captions across generated content.
Text-to-Video generation results. (Table 4,5)
Table 4: Overall performance of open-source Text-to-Video generation models over 10K Generate Any Scene captions.Red Cell is the highest score. Yellow Cell is the second highest score.Close-source models are evaluated on a 1K caption subset due to high inference cost.
Model clip score pick score
image reward
score
VQA score TiFA score
VideoCraft2 (Chen et al., 2024a) 0.2398 0.1976 -0.4202 0.5018 0.2466
[…]
ModelScope (Wang et al., 2023b) 0.2041 0.1886 -1.9172 0.3840 0.1219
Text2Video-Zero (Khachatryan et al., 2023) 0.2539 0.1933 -1.2050 0.4753 0.1952
CogVideoX-2B (Yang et al., 2024) 0.2038 0.1901 -1.2301 0.4585 0.1997
ZeroScope (Sterling, 2023) 0.2289 0.1933 -1.1599 0.4892 0.2388
[…]
Table 5: Overall performance of open-source Text-to-Video generation models over 10K Generate Any Scene captions with VBench metrics. Red Cell is the highest score. Blue Cell is the lowest score.
Model
subject
consistency
background
consistency
motion
smoothness
dynamic
degree
aesthetic
quality
imaging
quality
[…]
@@
[…]
@@
  1. 1.
    Open-source text-to-video models face challenges in balancing dynamics and consistency (Table 5). This is especially evident in Open-Sora 1.2, which achieves high consistency but minimal dynamics, and Text2Video-Zero, which excels in dynamics but suffers from frame inconsistency.
  2. 2.
  3. 3.
    As expected, SOTA close-source text-to-video models outperform others overall, particularly in image reward, VQA score, and TIFA score. This indicates their superior alignment with human preferences, as well as stronger faithfulness and compositional capabilities in generation.
  4. 4.
    Among open-source models, VideoCrafter2 strikes a balance across key metrics, leading in human-preference alignment, faithfulness, consistency, and dynamic.
Text-to-3D generation results. (Table 6)
@@
Table 6: Overall performance of Text-to-3D generation models over 1K Generate Any Scene captions. Evaluated on a 100 caption subset due to high inference cost.
Model clip score pick score vqa score tifa score
image reward
score
Latent-NeRF (Metzer et al., 2023) 0.2115 0.1910 0.4767 0.2216 -1.5311
DreamFusion-sd (Poole et al., 2022) 0.1961 0.1906 0.4421 0.1657 -1.5582
Magic3D-sd (Lin et al., 2023) 0.1947 0.1903 0.4193 0.1537 -1.6327
SJC (Wang et al., 2023a) 0.2191 0.1915 0.5015 0.2563 -1.4370
DreamFusion-IF (Poole et al., 2022) 0.1828 0.1857 0.3872 0.1416 -1.9353
Magic3D-IF (Lin et al., 2023) 0.1919 0.1866 0.4039 0.1537 -1.8465
ProlificDreamer (Wang et al., 2024c) 0.2125 0.1940 0.5411 0.2704 -1.2774
Meshy-4 (Meshy AI, 2025) 0.2163 0.1922 0.5290 0.2908 -1.0496
  1. 1.
    Among open-source models, ProlificDreamer outperforms other models, particularly in ImageReward Score, VQA Score and TIFA Score.
  2. 2.
  3. 3.
    Meshy-4 demonstrates overall superior performance compared to all open-source models, especially in terms of Clip Score, TIFA Score and ImageReward Score, reflecting its strengths in semantic generation and human preference alignment.
A.3 Validation of Phrasing Robustness and Human Alignment
To assess robustness to linguistic variation and to verify that automated metrics reflect human preferences, we conduct two focused studies.
A.3.1 Phrasing Robustness via Paraphrasing
Setup. We sample 100 scene graphs from the 10K benchmark while preserving the distribution of object counts, relation density, and attribute complexity. For each graph, GPT-4o generates a linguistically varied yet graph-faithful caption using the prompt below.
Paraphrasing Prompt You are given a scene graph in JSON format, where:
-- "nodes" contain objects and their attributes,
-- "edges" describe relationships between objects or link attributes to objects.
Your task:
1. Understand the semantic meaning of each node and edge.
2. Convert the graph into a natural language caption that describes the entire scene.
3. Include all objects, attributes, and relations from the graph, and strictly follow the graph structure.
4. Do not introduce new objects or relationships not present in the graph.
Input: {scene_graph}
We then re-score all models with VQA Score under these paraphrased captions. Results are listed in Table 7.
Table 7: Paraphrase robustness: VQA Score and ranks on 100 graphs.
Model Orig. Score Para. Score Diff Orig. Rank Para. Rank
DALLE-3 0.6871 0.7542 +0.0671 1 1
FLUX.1-schnell 0.6132 0.6648 +0.0516 2 2
PixArt-$\Sigma$ 0.6109 0.6159 +0.0050 3 3
PixArt-$\alpha$ 0.6049 0.6043 -0.0006 4 4
Playground v2.5 0.5734 0.5075 -0.0659 5 8
Stable Diffusion 3 0.5579 0.5140 -0.0439 6 7
FLUX.1-dev 0.5561 0.5024 -0.0537 7 9
DeepFloyd IF XL 0.5397 0.5606 +0.0209 8 5
Wuerstchen v2 0.5352 0.5014 -0.0338 9 10
SDXL 0.5328 0.5322 -0.0006 10 6
SD v2-1 0.5282 0.4961 -0.0321 11 11
Findings.
The Pearson correlation coefficient between model rankings on programmatic versus paraphrased captions is 0.9232, indicating a very strong positive correlation.
This validation study demonstrates strong consistency between the two approaches. Importantly, the top-performing models (DaLL-E 3, FLUX.1-schnell, $\Sigma, PixArt$-$\alpha$) maintain their rankings across both evaluation conditions, while the relative ordering of models remains largely consistent. This high correlation validates that our programmatic approach produces rankings that are generalizable and not artifacts of the templated caption generation. The slight variations observed (e.g., some mid-tier models showing small rank changes) are within expected bounds and do not affect the overall conclusions about model capabilities.
A.3.2 Human Alignment Study
Setup. We evaluate six representative models (DaLL-E 3, FLUX.1-schnell, PixArt-$\Sigma$, Playground v2.5, SD3 Medium, SDv2.1) with diverse performance characteristics and recruit 3 human evaluators. Three independent evaluators each assess 40 caption–image groups, with 10 shared overlapping groups across all evaluators to measure inter-annotator agreement. Evaluators ranked the generated images based on both relevance to the caption and overall visual quality. We show the rankings in Table 8.
Findings
Table 8: Human vs. VQA rankings (lower is better).
Model VQA Rank Human Avg. Rank
DaLL-E 3 1 1
FLUX.1-schnell 2 2
PixArt-$\Sigma$ 3 4
Playground v2.5 4 3
SD3 Medium 5 5
SDv2.1 6 6
Inter-annotator reliability. The 3 evaluators showed strong agreement on the 10 shared samples, with a Spearman correlation coefficient of 0.962, demonstrating consistent human judgment criteria.
Human–metric alignment. The correlation between human rankings and our VQA Score rankings is 0.918, indicating strong alignment between automated and human evaluation:
This study validates that our VQA Score-based rankings closely align with human preferences. The consistency between automated metrics and human judgment strengthens confidence in our benchmark’s ability to assess model performance in a manner that reflects human perception.
A.4 More Analysis with Generate Any Scene
With Generate Any Scene, we can generate infinitely diverse and highly controllable captions. Using Generate Any Scene, we conduct several analyses to provide insights into the performance of today’s Text-to-Vision generation models.
A.4.1 Performance analysis across caption properties
Perplexity. (Figure 8)
Commonsense. (Figure 9)
Commonsense is an inherent property of text. We utilize the Vera Score (Liu et al., 2023), a metric generated by a fine-tuned LLM to evaluate the text’s commonsense level.
Element Numbers (Complexity of Scene Graph). (Figure 10)
A.4.2 Analysis on different metrics
Clip Score isn’t a universal metric.
Clip Score is one of the most widely used metrics in Text-to-Vision generation for evaluating the alignment between visual content and text. However, our analysis reveals that Clip Score is not a perfect metric and displays some unusual trends. For instance, as shown in Figures 8, 9, and 10, we compute the perplexity across 10K captions used in our study, where higher perplexity indicates more unpredictable or disorganized text. Interestingly, unlike other metrics, Clip Score decreases as perplexity lowers, suggesting that Clip Score tends to favor more disorganized text. This behavior is counterintuitive and highlights the potential limitations of using Clip Score as a robust alignment metric.
Limitations of human preference-based metrics.
We use two metrics fine-tuned using human preference data: Pick Score and ImageReward Score. However, we found that these metrics exhibit a strong bias toward the data on which they were fine-tuned. For instance, as shown in Table 3, Pick Score assigns similar scores across all models, failing to provide significant differentiation or meaningful insights into model performance. In contrast, ImageReward Score demonstrates clearer preferences, favoring models such as DaLL-E 3 and Playground v2.5, which incorporated human-alignment techniques during their training. However, this metric shows a significant drawback: it assigns disproportionately large negative scores to models like SDv2.1, indicating a potential over-sensitivity to alignment mismatches. Such behavior highlights the limitations of these metrics in providing fair and unbiased evaluations across diverse model architectures.
VQA Score and TIFA Score are relative reliable metrics.
A.4.3 Fairness analysis
Gender.
Figure 11: Average performance scores of all models across different genders evaluated using various metrics.
Race.
There are also performance gaps in different races. From Figure 12, we found that "white (person)" and "black (person)" perform better than "asian (person)", "Indian (amerindian)", and "Latin American".
A.4.4 Correlation of Generate Any Scene with other Text-to-Vision generation benchmarks
The Generate Any Scene benchmark uniquely relies entirely on synthetic captions to evaluate models. To assess the transferability of these synthetic captions, we analyzed the consistency in model rankings across different benchmarks (Wu et al., 2024b; Li et al., 2024a; Wu et al., 2023). Specifically, we identified the overlap of models evaluated by two benchmarks and computed the Spearman correlation coefficient between their rankings.
As shown in the figure 13, Generate Any Scene demonstrates a strong correlation with other benchmarks, such as Conceptmix (Wu et al., 2024b) and GenAI Bench (Li et al., 2024a), indicating the robustness and reliability of Generate Any Scene’s synthetic caption-based evaluations. This suggests that the synthetic captions generated by Generate Any Scene can effectively reflect model performance trends, aligning closely with those observed in benchmarks using real-world captions or alternative evaluation methods.
(a) SDv2.1 vs. SD3 Medium on average VQA Score in fine-grained categories.
(b) PixArt-$\Sigma$ vs. SD3 Medium on average VQA Score in fine-grained categories.
(c) FLUX.1-schnell vs. SD3 Medium on average VQA Score in fine-grained categories.
(d) PixArt-$\Sigma$ vs. FLUX.1-schnell on average VQA Score in fine-grained categories.
Figure 14: Pairwise comparison on average VQA Score in fine-grained categories.
A.4.5 Case study: Pairwise fine-grained model comparison